OpenStudy (babystar16):

if 8 pens and 7 pencils cost $3.37 while 5 pens and 11 pencils cost $3.10, how much does each pen and pencil cost?

OpenStudy (mathlegend):

@babystar16 do you want me to show you how to solve?

OpenStudy (babystar16):

plzzzzzzzzzzzzzzz

OpenStudy (mathlegend):

So we can setup system of linear equation.

OpenStudy (mathlegend):

P = Pens C = Pencils

OpenStudy (mathlegend):

8p+7c=3.37 5p+11c=3.10

OpenStudy (mathlegend):

@babystar16 can't we multiply both of the 'p' variables to get a common number? Lets multiply the 8p times -5. Then the 5p times 8.

OpenStudy (mathlegend):

-40p-35c=-16.85 40p+88c=24.80 ------------------ (Now, we add...) 53c = 7.95 c = .15

OpenStudy (mathlegend):

@babystar16 do you understand this?

OpenStudy (mathlegend):

Now, just plug back in .15 for c in the original equation.

OpenStudy (mathlegend):

5p+11(.15)=3.10 5p + 1.65 = 3.10 5p = 1.45 p = .29

OpenStudy (mathlegend):

Pens were .29 cent Pencils were .15 cent

OpenStudy (babystar16):

thank u

OpenStudy (babystar16):

thank u for helping me understand it